Grade upto college level Physical Chemistry

For the reaction, N2O5(g) = 2NO2(g) + 0.5 O2(g), calculate the mole fraction of N2O5(g) decomposed at constant volume temperature, if the initial pressure is 600 mm Hg and the pressure at any time is 960 mm Hg. Assume ideal gas behavior.

Shane Macguire
12 Years agoGrade upto college level

1 Answer

Navjyot Kalra
12 Years ago
N2O5(g) ⇌ 2NO2(g) + 1/2 O2(g)
Initial pressure 600 0 0
Final pressure 600 – P 2P P/2
P ∝ moles when V and T are constant
(where moles equivalent to pressure –P are decomposed)
Total pressure = 600 – P + 2 P + P/2 = 960 mm of Hg
∴ P = 240 mm Hg
Thus moles of N2O5 decomposed = 240/600 = 0.4